OpenStudy (anonymous):

simplify 4x^2+12x-16/2x+10 divided by 6x+24/x^2+9x+20

OpenStudy (saifoo.khan):

\[\frac{4(x-1)(x+4)}{2(x+5)} \times \frac{(x+4)(x+5)}{6(x+4)}\]Can u solve now?